### v1

# ZeroCAP: Zero-Shot Multi-Robot Context Aware Pattern Formation via Large Language Models

Abstract
Incorporating language comprehension into robotic operations unlocks significant advancements in robotics, but also presents distinct challenges, particularly in executing spatially oriented tasks like pattern formation. This paper introduces ZeroCAP, a novel system that integrates large language models with multi-robot systems for zero-shot context aware pattern formation. Grounded in the principles of language-conditioned robotics, ZeroCAP leverages the interpretative power of language models to translate natural language instructions into actionable robotic configurations. This approach combines the synergy of vision-language models, cutting-edge segmentation techniques and shape descriptors, enabling the realization of complex, context-driven pattern formations in the realm of multi robot coordination. Through extensive experiments, we demonstrate the systems proficiency in executing complex context aware pattern formations across a spectrum of tasks, from surrounding and caging objects to infilling regions. This not only validates the system’s capability to interpret and implement intricate context-driven tasks but also underscores its adaptability and effectiveness across varied environments and scenarios.
I Introduction
Pattern formation, a key idea in the organization of multi-robot systems [1], involves arranging robots in a specific way to reach a common objective. This skill is essential for various purposes such as environmental surveillance, precise farming, cooperative transportation, and emergency management. The significance of pattern formation lies in its capacity to utilize the combined actions of robots to carry out tasks that go beyond what individual robots can achieve, thus improving effectiveness, adaptability, and robustness. Central to this endeavor is not just the geometric arrangement of robots, but the evolution towards purposeful pattern formations—where the design and execution of patterns are intrinsically linked to the context and objects within the environment. This nuanced approach marks a significant departure from traditional methodologies, steering the field towards more interactive and object-centric tasks that can be applicable to real world scenarios as shown in Fig 1.
Despite its significance, traditional approaches to pattern formation in robotics have been predominantly constrained by model/rule-based methods and machine learning paradigms. These methodologies, while foundational, have their limitations: model rule-based strategies [1, 2, 3, 4], are often tailored to specific scenarios/patterns and lack the flexibility to adapt to new or unmodeled patterns. Reinforcement learning (RL) approaches, on the other hand, require training time and extensive reward design, making them impractical for scenarios requiring immediate deployment. Moreover, machine learning (ML) based methods [5, 6, 7] demand large datasets of annotated or labeled data for training, a requirement that poses substantial challenges in dynamic or novel environments. These limitations across all the methodologies are particularly pronounced in scenarios where pattern formations must be conditioned with respect to specific objects of interest, underscoring the need for a more adaptable and intuitive methodology.
Figure 1: A conceptual image illustrating the proposed Zero-Shot multi-robot Context Aware Pattern (ZeroCAP) formation system, beginning with an initial environment where a car is parked incorrectly. Upon receiving the command from human operator to “Surround the incorrectly parked car at the corners!”, the system identifies the target vehicle and autonomously positions the robot cones at strategic corner locations, showcasing the successful execution of a zero-shot context aware pattern formation task.
Integrating language comprehension into robotic systems heralds considerable progress [8, 9], yet it poses unique challenges, especially in spatially-focused tasks such as pattern formation. While language models have shown a remarkable ability as general pattern machines [10] and offer a robust framework for parsing and generating human-like text, Vision Language Models (VLMs) falter in accurately localizing spatial tasks, accentuating the need for techniques that adeptly marry abstract instruction understanding with their tangible implementation.
Our contribution to this evolving field addresses the challenge previously identified by [3], focusing on the dynamic assignment of a variable number of robotic agents to specific target positions for intricate pattern formations. By combining vision-based processing with the sophisticated interpretative capabilities of Large Language Models (LLMs), we introduce a Zero-Shot multi-robot Context Aware Pattern (ZeroCAP) formation system that facilitates zero-shot, language, and context-conditioned pattern formation. It processes both images and natural language instructions to identify objects of interest and generate precise coordinates for robot deployment, enabling immediate and purposeful pattern formation directly influenced by user instructions.
We distinguish our approach by categorizing pattern formations into three purposeful types: 1) General pattern formation, where robots arrange themselves around the object of interest based on user specification; 2) Infill pattern formation, where robots occupy the interior space of a region based on user specification; and 3) Caging pattern formation, creating a boundary around the object. This categorization increases the flexibility of deployment and signifies a shift from standard geometric patterns to contextually relevant formations tailored to the specific user context. Introducing purposeful pattern formation, our system adapts to environmental cues through natural language instructions, demonstrating language and context-driven pattern formation’s viability.
The following sections will explore the technical challenges addressed, methodologies applied, and the significant impact our work has on advancing robotics towards intuitive, flexible, and context-aware multi-robot systems.
II Related Works
Pattern Formation. Multi-robot pattern formation research has explored centralized control [11], [12] and distributed approaches [1], [13]. Centralized systems face scalability and adaptability challenges, while distributed methods, including consensus-based [14], leader-follower [15], and bio-inspired [16], improve scalability but lack robustness due to pattern-specific, rule-based strategies. Advancements to the feild include potential field-based methods[17], [4] and artistic formations [3], plus novel algorithms such as event-triggered mechanisms [2]. However, model/rule-based strategies typically are designed for specific scenarios, lacking the adaptability required for new or unseen patterns or even handling user specific instructions. Conversely, learning-based strategies like Q-learning [7], [6]while offering more generalizability are often time intensive with their training and are limited to the fixed patterns that the system was trained for. Despite the above progress over the years, significant challenges persist in creating flexible solutions that quickly adapt to contextual user preferences and situations without heavy pre-configuration, training, or extensive information exchange. These issues highlight a broader need in the feild for moving from traditional geometric patterns to contextually relevant formations tailored to context specific user tasks and environments.
Our approach diverges from these traditional paradigms by adopting zero-shot learning using a language model for robotics, sidestepping the limitations of both rule-based and machine learning methodologies. By leveraging natural language instructions, our system achieves a flexible, intuitive means of directing multi-robot pattern formation, capable of adaptation to novel scenarios without the need for extensive dataset training or predefined rules. This represents an advancement over existing methods centralized or decentralized.
LLM for Robotics. LLMs are pivotal in robotics for task planning through zero-shot learning [18, 19, 10], excelling in diverse plan generation [20, 9, 8] and performance refinement [21]. In the realm of multi-agent systems, LLMs have been instrumental in fostering agent cooperation, equipping individual agents with the capabilities to improve interaction dynamics and collaborative efficacy [22, 23, 24]. Despite their growing utility in robotics, none of these noteable methodologies cover pattern formation among multi-robot systems. Vision Language Models, while useful in many applications involving robot navigation[25] and high-level action control[26, 27], are limited by their spatial reasoning capabilities [28], impacting their effectiveness in robotics applications that require precise spatial coordination. Our research leverages VLMs and LLMs for multi-robot pattern formation, addressing the spatial task limitations of VLMs and pioneering a linguistically-driven approach to complex robotic coordination aided by representing shapes using a node graph.
III Problem Description
Our research delves into the nuanced task of deploying robotic agents to form patterns around or inside a designated object of interest $O$, located within an environment, $e$. This environment is depicted through an image ($I_{E}$), paired with a natural language instruction ($Ins$) that outlines the desired pattern formation task. Operating under a centralized control system, robots, simplified to points of mass, are coordinated to ensure precise pattern formation in the context of $O$ and $Ins$.
The procedure begins with the identification and segmentation of $O$ from $I_{E}$, leading to a mask image that segregates $O$ from the rest of $e$. This step is crucial to accurately define the geometric shape of $O$. The shape of $O$ is mathematically modeled as a connected graph $G=(V,E)$ [29], where $V$ signifies the set of vertices or nodes, outlining discrete points that mark the shape’s corners or distinct features, and $E$ represents the set of edges connecting these vertices, thus framing the pattern without self-connectivity.
Let $R=\{r_{1},r_{2},\ldots,r_{N}\}$ represent the list of robotic agents available, where $N$ is the total number of robots. Given $Ins$, and $I_{E}$, the system, through a series of steps that include object detection, segmentation, and shape determination resulting in $G$, leverages an LLM to interpret and reason natural language instructions in light of $G$. The output of the system is a list of $(x,y)$ coordinates, $C_{xy}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{N},y_{N})\}$, which assigns each robot $r_{i}$ in $R$ to a specific coordinate $(x_{i},y_{i})$ for deployment. This assignment facilitates the formation of the specified pattern around/inside $O$, in accordance with $Ins$.
Formally, this process can be encapsulated as a function $F$, where $F(I_{E},Ins)\rightarrow C_{xy}$, mapping the inputs to the deployment coordinates for each robot.
Figure 2: An overview of the ZeroCAP system. It traces the workflow from the initial natural language instruction and input image of the environment through to the final deployment of robots, illustrating the sequence of processing stages—including context identification using Vision Language Model (VLM), object segmentation, shape description, and Large Language Model (LLM) coordination for precise robot placement in the environment. Three key stages are highlighted and explained in Section IV.
IV ZeroCAP Formation System
This section delineates the comprehensive methodology of our ZeroCAP formation system, which introduces a novel approach for zero-shot multi-robot context aware pattern formation. Leveraging advanced VLMs, the methodology integrates linguistic instructions with visual perception, enabling precise spatial reasoning and object identification in multi-robot systems. Our approach harnesses the synergy of cutting-edge segmentation techniques, and utilizes standard computer vision algorithms for obtaining shape description. The decision-making process is empowered by a specialized LLM, which synthesizes the spatial description of object and linguistic context to generate target deployment coordinates for the robotic agents. The multi-robot control is centrally orchestrated, ensuring efficient navigation. The following subsections provide a detailed exposition of each stage, from initial input processing to the final deployment of robots.The overall structure of the proposed ZeroCAP framework is shown in Fig. 2.
IV-A Context Identification using VLM
The initiation of our methodology engages a VLM to interpret an environmental image $I_{E}$ and a natural language instruction $Ins$. The architecture of this model excels at correlating visual stimuli with corresponding linguistic descriptors, emulating human-like proficiency in image-language tasks. Provided with $I_{E}$ and $Ins$, the VLM processes these inputs in tandem, producing two key outputs: the Object of Interest $O$ and the Pattern Formation Instruction $PI$.
The Object of Interest $O$ is identified utilizing the VLM’s zero-shot detection capabilities, circumventing the need for prior object-specific training and ensuring versatility within varying environments. Conventional object detection methods may localize objects but lack the sophisticated reasoning to discern context as VLMs do. Given an instruction akin to that of Fig. 3, standard natural language-aided object detectors [5] would falter in reasoning with which object is odd. In contrast, the advanced reasoning capabilities of the VLM, contextualized by the instruction, render it superior for identifying objects of interest within complex scenarios. This nuanced understanding of context allows the VLM not only to detect but also to reason about objects in a manner aligned with the given linguistic directives. Concurrently, the Pattern Formation Instruction $PI$ is extrapolated from the natural language directive enriched by contextual nuances in $I_{E}$. Here, the VLM generates $PI$, inspired by chain-of-thought reasoning [30], to discern spatial and context-sensitive task directives. In Algorithm 1, Line 1 corresponds to estimating the context using the VLM.
One notable strength of our system is our approach to mitigating the common issue of precise spatial localization, a challenge in standard VLMs. The inherent limitations of VLMs in tasks that require exact positional outputs make it impractical to use these models to directly provide the deployment coordinates $C_{xy}$. Instead, the VLM is optimized to identify $O$ with necessary precision and to extract a context-rich $PI$. These are indispensable for subsequent segmentation and precise pattern formation stages conducted by more specialized subsystems. By honing in on general features and attributes discernible in the environment and the language input, the VLM identifies $O$ with the necessary precision for subsequent segmentation and pattern formation stages.
The VLM’s output, consisting of $O$ and $PI$, thus propels the ZeroCAP system into the segmentation phase, where $O$ is visually distinguished within $I_{E}$, and $PI$ informs the strategic pattern formation instruction to be implemented by the robotic entities.
Figure 3: An illustration of the VLM processing a natural language instruction to identify the object of interest and generate a pattern formation instruction, which guides the deployment of robots in a specified arrangement around the object. The object is not explicitly mentioned in the instruction and must be reasoned by the VLM within the context of image of the environment.
IV-B Object Segmentation and Shape Description
Following the identification of $O$ and the acquisition of the $PI$ from the VLM, the system transitions into the crucial stages of segmentation and shape description. These stages leverage sophisticated language conditioned segmentation techniques [5, 31], to perform image segmentation, followed by shape description stage using standard computer vision methods.
IV-B1 Object Segmentation
The segmentation process begins with language-prompt-based object segmentation, which employs LangSAM, an open-source language-conditioned segmentation model built upon segment-anything approach and the GroundingDINO detection model [31, 32]. Provided with the Object of Interest $O$ and the Input Image $I_{E}$, these models generate a binary segmentation map in which the object is highlighted in white against a black background.
Upon completion, the segmentation yields a Masked Image $I_{M}$ and an accuracy metric $\alpha$, compared against a predefined threshold $t$. If $\alpha$ meets or exceeds this threshold, the process proceeds, confirming the segmentation’s quality for subsequent tasks. This segmentation approach ensures the precise extraction of $O$ from $I_{E}$, encapsulated within $I_{M}$ for the next processing stage. Lines 2 to 4 of the Algorithm 1 expound the segmentation process.
IV-B2 Shape Description
With a high quality $I_{M}$ at hand, the ZeroCAP system advances to the shape description phase. Here, standard computer vision techniques preprocess $I_{M}$, including filtering and enhancement operations to sharpen feature delineation. Through OpenCV’s edge detection algorithms and contour detection methods, $I_{M}$ is analyzed to determine $O$ geometric form. This analysis yields the object’s edges $E$ and vertices $V$, forming a connected graph $G=(V,E)$ that mathematically represents the object’s structure. Algorithm 1 delineates this process in Lines 10 – 13, with a method call in Line 5.
Integrating these segmentation techniques with precise shape description, the ZeroCAP system ensures accurate recognition, segmentation, and modeling of the Object of Interest. This forms a solid foundation for the ensuing methodology phase: LLM-based Pattern Former.
Algorithm 1 ZeroCAP System Pipeline
1:Image of the Environment $I_{E}$, Natural Language Instruction Ins
2:List of coordinates for robot deployment $C_{xy}$
3:$(O,PI)\leftarrow\textsc{VLM.Process}(I_{E},\textit{Ins})$
4:$(I_{M},\alpha)\leftarrow\textsc{Segment}(O,I_{E})$
5:if $\alpha<t$ then
6:   return “Insufficient segmentation accuracy.”
7:$(V,E)\leftarrow\textsc{ShapeDescriptor}(I_{M})$
8:$C_{xy}\leftarrow\textsc{LLM.GenerateCoordinates}(PI,V,E,R)$
9:for each coordinate $(x_{i},y_{i})$ in $C_{xy}$ do
10:   $r_{i}\leftarrow R[i]$
11:   DeployRobots($r_{i},(x_{i},y_{i})$)
12:procedure ShapeDescriptor($I_{M}$)
13:   Preprocess the masked image $I_{M}$
14:   Extract edges $E$ and vertices $V$
15:   return $(V,E)$ from $I_{M}$
16:procedure DeployRobots($r,(x,y)$)
17:   Command robot $r$ to move to coordinates $(x,y)$
18:   Verify that robot $r$ has reached $(x,y)$
IV-C LLM-based Pattern Former
Upon accurately delineating the geometry of $O$, the context-aware multi-robot deployment system advances to determining goal position coordinates using a LLM. This stage involves the LLM being provided with the number of available robots $R$, interpreting the Pattern Formation Instruction $PI$ alongside the object’s spatial characteristics, represented through edges $E$ and vertices $V$ to provide an output representing positions for robots.
The LLM’s task is to integrate the $PI$ with the spatial data of $E$ and $V$ to determine the optimal pattern formation. This involves comprehensive reasoning to align the robots according to $PI$, while considering the shape of the object and the available robotic agents $R$.
In particular, our methodology eschews the direct use of the entire Masked Image $I_{m}$ as input to the LLM, a decision underscored by considerations of computational efficiency and token costs. Direct image is resource-intensive, whereas leveraging $E$ and $V$ markedly reduces computational load especially when accounting for tokens in the context of LLM’s. This choice highlights our system’s capacity for real-time processing and deployment efficiency while also remaining inexpensive in computation and practicality.
The culmination of this process is the production of a coordinate list $C_{xy}$ as seen in Line 6 of Algorithm 1, which prescribes exact deployment coordinates for each robot.
IV-D Robot Deployment
Upon generating the deployment coordinates $C_{xy}$ with the LLM, the ZeroCAP system initiates robot deployment, translating algorithmic output into coordinated physical actions.
Each robot $r_{i},i\in N$ is directed toward its assigned location $(x_{i},y_{i})$ under centralized control, ensuring precise navigation. Upon arrival, the system verifies each robot’s position against $C_{xy}$, adjusting as necessary to rectify any positional inaccuracies, thereby ensuring the formation aligns perfectly with the pattern instructions. This is delineated in Lines 7 to 9 and 14 to 16 in Algorithm 1.
The successful deployment of robots in the context of the Object of Interest $O$ marks the completion of the pattern formation task, embodying the ZeroCAP system’s capability to execute context-driven, instruction-based patterns through autonomous robotic agents.
V Experiment
(a) General pattern formation tasks: robots arrange themselves around the object of interest based on user specification
(b) Infill pattern formation tasks: robots occupy the interior space of a region based on user specification
(c) Caging formation tasks: creating a cage like boundary around the object
(d) Application oriented task: surrounding the incorrectly parked car with four robots.
Figure 4: Illustrations of the three types of pattern formation tasks and a task demonstration of Fig 1, showcasing the system’s adaptability to various instruction types and object configurations. In all tasks, robots are represented as black points initially distributed at the bottom of the scene in the initial environment, transitioning to form user-directed patterns in the final environment. This illustrates the dynamic capability of our system to interpret instructions and spatially reorganize the robots accordingly. Please zoom in on the images for details.
Our experimental evaluation investigates the ZeroCAP system’s zero-shot pattern formation capabilities across ten tasks as shown in Fig. 4, delineated into three categories as prefaced in Section I: general pattern formations as shown in subfigure 4(a), infill pattern formations as shown in subfigure 4(b), and caging formation tasks as shown in subfigure 4(c) and the task referred to in Fig 1 as shown in subfigure 4(d).
Tasks are diversified by object setup to evaluate the system’s adaptability and discernment capabilities under varying environmental complexities. In the Single Object setup (ENV a, b, d, g), the environment contains a solitary object, simplifying the identification process. The Multi-Object setup (ENV c, e, h) introduces additional objects as distractors, challenging the system to isolate the target object based on explicit user instructions. Lastly, the Hidden Object setup (ENV f, i, park) presents a more complex scenario where the natural language instruction implies the object of interest without explicit mention, compelling the system to employ reasoning to deduce the correct object.
V-A Baseline Comparisons
Our system’s performance was benchmarked against three baselines. The first is an RL-based simplified approach with hand-crafted rewards, serving primarily as a pattern-formation benchmark without natural language comprehension. The second and third baselines use GPT4 Vision [33] and llava-v1.6 [34], respectively, relying solely on vision-language models to complete the tasks. The LLMs employed in our experiments are GPT-4 [35], llama70b [36], and Claude [37], contrasting their performance against each other and the baselines. For ablation studies, the default configuration of the ZeroCAP system utilizes GPT-4 as the primary LLM.
V-B Environment
Experiments were conducted in a 2D simulation using Pygame, with robots as point-mass entities capable of precise movements to target coordinates. The setup includes multi-object tasks with infill and caging requirements, omitting obstacle avoidance behaviors to focus on the pattern formation tasks. A subset of tasks was also executed on a real testbed, featuring Hamster mobile robots [38] under an overhead camera setup. Fig. 5 showcases two tasks executed in a real-world setting, serving as a tangible proof of concept that corroborates our simulation results. To facilitate these tasks, the environment is simplified into a grid-based image, with each grid box approximately matching the size of a robot. This adaptation allows the robots to retain their point mass entity characterization within the practical constraints of the real-world experiments, ensuring consistency in their deployment and movement. For demonstrations of simulator and real-world tasks, please refer to the supplementary video.
V-C Evaluation Metrics
We employed two main evaluation metrics: Success Rates (SR) and Goal Condition Recall (GCR). SR indicates whether the robots formed the pattern successfully, verified by handcrafted ground truth labels, while GCR quantifies the proportion of robots correctly positioned within the pattern. A GCR of 1 implies a perfect match with the ground truth and, consequently, an SR of 1. Each task underwent 10 trials, with averages reported for SR and GCR, chosen to account for the inherent variability in LLM performance.
VI Results and Analysis
TABLE I: Comparative analysis of pattern formation methodologies across General, Infill, and Caging tasks, evaluating Success Rates (SR) and Goal Condition Recall (GCR) metrics across different methods.
Methods General Infill Caging Total
SR GCR SR GCR SR GCR SR GCR
GPT-4 Vision (Only) 0 0 0 0 0 0 0 0
LLaVA v1.6 (Only) 0 0 0.13 0.1 0 0.07 0 0.07
RL Baseline 0.35 0.53 0.28 0.41 0.88 0.9 0.5 0.61
ZeroCAP (GPT-4) 0.71 0.85 0.86 0.89 0.74 0.82 0.77 0.85
ZeroCAP (lama-2-70b) 0.2 0.47 0.39 0.71 0.32 0.67 0.3 0.61
ZeroCAP (claude-3-opus) 0.57 0.78 0.72 0.88 0.64 0.73 0.64 0.79
Table I shows a comprehensive evaluation of different methodologies tested in this paper, revealing the effectiveness of the ZeroCAP system in comparison to baselines. The initial baselines, GPT-4 Vision and LLava v1.6, were employed to evaluate the feasibility of using VLMs directly for zero-shot context-aware pattern formation. The performance of these baselines was significantly poor. GPT-4 Vision generally failed to generate any coordinates, reflecting its known limitations with spatial localization tasks. LLava, while capable of producing coordinates, generated positions that were largely inaccurate and contextually irrelevant. These outcomes effectively demonstrate that direct application of VLMs for context-aware pattern formation is non-trivial, and without additional processing or reasoning mechanisms, VLMs fall short of providing viable solutions. The RL baseline exhibited a mixed performance across the task spectrum. It notably struggled with the general pattern formation tasks, where user specifications varied greatly and required a high degree of interpretation and flexibility. The RL method’s rigidity and lack of language comprehension rendered it unable to adapt to the nuanced demands of these tasks. On the other hand, the RL approach showed considerable success in Caging tasks, aligning with the more straightforward objective of these tasks where the overarching goal is consistently to form a barrier around the object.
Figure 5: Two real-world scenarios demonstrating ZeroCAP system’s ability to execute context-aware pattern formations.
The ZeroCAP methods, incorporating LLMs like GPT-4, llama-2-70b, and claude-3-opus, present a contrasting picture. The GPT-4 based ZeroCAP achieves the highest overall SR and GCR across all types of tasks, indicating its superior pattern recognition and execution in zero-shot conditions. This suggests that GPT-4’s advanced natural language processing abilities are effectively leveraged by the ZeroCAP system to interpret instructions and generate appropriate robotic behaviors. The performance of ZeroCAP using llama-2-70b and claude-3-opus varies. While the llama-2-70b variant has some success, its inconsistency across tasks points to possible limitations in its ability to handle complex spatial reasoning as compared to GPT-4. Claude-3-opus shows a relatively more consistent performance. Across different pattern formation types, ZeroCAP methods behave in a manner that suggests certain patterns may be more conducive to LLM-driven execution. The General pattern formation tasks appear to be the most challenging, possibly due to the abstract nature of the instructions and the need for more nuanced interpretation. Infill tasks, which involve a more defined area and clear objectives being in most cases to deploy robots inside a region, show the highest SR and GCR for GPT-4, indicating a strong alignment between the method’s capabilities and the task requirements. Caging tasks, while more structured, still present a challenge, but ZeroCAP methods demonstrate reasonable effectiveness.
VI-A Ablation: Segmentation Methods
The ablation study presented in Table II examines the efficacy of CLIPSeg [5] and LangSAM as object segmentation methods within the context of language-prompted object identification. These methods, while capable of segmenting objects based on language instructions, do not possess the same reasoning abilities provided by full-fledged VLMs. The table compares the performance of these segmenters used in isolation (denoted by ‘Only’) to their performance when integrated into the ZeroCAP framework.
TABLE II: Ablation studies comparing the success rates (SR) of standalone segmentation methods and their integration within the ZeroCAP system.
Method SR
CLIPSeg (Only) 0.51
LangSAM (Only) 0.43
ZeroCAP with CLIPSeg 0.78
ZeroCAP with LangSAM 0.77
It is evident that using CLIPSeg and LangSAM alone for object identification and segmentation results in modest success rates of 0.51 and 0.43, respectively. This suggests that while they can interpret natural language to some extent and perform segmentation, the lack of deeper contextual understanding and spatial reasoning affects their overall effectiveness in pattern formation tasks. This is especially a factor in case of hidden object task scenarios.
When these segmentation methods are incorporated into the ZeroCAP system, the success rates improve significantly to 0.78 for CLIPSeg and 0.77 for LangSAM. The integration of these segmenters within ZeroCAP likely benefits from additional contextual processing and spatial reasoning capabilities provided by the LLMs in the system, leading to more successful task outcomes.
VI-B Ablation: Shape Descriptors
The ablation study presented in Table III evaluates the influence of different shape descriptors on the performance of the ZeroCAP system. ‘Increase in Tokens’ indicates the relative percentage increase in the number of tokens used by the LLM compared to using only ‘Edges’ as the shape descriptor, which serves as the baseline in this context.
TABLE III: Ablation studies on the impact of different shape descriptors on Success Rates (SR) and their relative increase in tokens required for LLM processing.
Shape Descriptor SR
Increase
in Tokens
Edges 0.77 -
Edges and Vertices 0.48 2%
Binary Image Matrix 0.71 400%
Analyzing the data, it is evident that using only ‘Edges’ as a shape descriptor achieves the highest SR of 0.77. This suggests that ‘Edges’ alone provide ample information for the LLM to accurately generate the desired patterns without requiring additional token input. In contrast, when ‘Edges and Vertices’ are combined, there is a modest increase in tokens by 2%, but the SR significantly decreases to 0.48. This decrease could imply that including vertices introduces a level of complexity that the LLM does not effectively translate into better pattern formation. Moreover, vertices as a descriptor could potentially be misleading for non-polygonal shapes where the definition of vertices is less clear, potentially leading to inaccurate pattern formation by the LLM.
The binary image matrix descriptor is obtained by converting the segmented object mask into a binary matrix, which is then resized to a 100x100 grid to fit within the token constraints of the LLM. This matrix serves as a input for the LLM, representing the shape in a form that can be directly processed. The LLM utilizes this grid-based representation to understand the spatial arrangement and generate target coordinates for the robots. The binary image matrix descriptor, incurs a 400% increase in token usage. This upsurge poses practicality concerns due to token limits and practical costs. Additionally, resizing may obscure vital shape details, affecting pattern accuracy. Despite a reasonable SR of 0.71, the trade-off between detail richness and token efficiency is crucial for the ZeroCAP system’s viability.
VII Discussion
The research presented in this paper marks a significant stride in the domain of multi-robot systems, particularly in the application of LLMs for zero-shot context aware pattern formation. Nonetheless, the current scope is confined to static patterns, leaving the exploration of dynamic patterns - those that evolve over time in response to environmental changes or ongoing tasks - as a promising avenue for future research. Dynamic pattern formation would not only extend the flexibility and adaptability of multi-robot systems but also increase their applicability in real-world scenarios such as search and rescue operations or environmental monitoring, where conditions are continually shifting.
Our current method’s reliance on centralized processing poses potential bottlenecks and failure points in scenarios requiring quick robot adaptability. Exploring decentralized or hybrid control systems could improve robustness and scalability by merging centralized planning with decentralized execution. Additionally, incorporating sensory feedback mechanisms may enhance decision-making, maintaining system effectiveness amid real-time environmental changes.
Our research lays a foundation for future advancements in robotic pattern formation and practical applications, extending from complex coordination tasks in surveillance, logistics, and warehousing to context-aware strategies. Progressing beyond current limitations and enhancing our foundational framework is vital for fully realizing robotics’ capabilities in intricate, context-sensitive environments.
VIII Conclusion
Our research has successfully introduced ZeroCAP, an innovative framework that harmonizes the advanced linguistic comprehension of large language models with the physical capabilities of robotic systems to achieve zero-shot pattern formation. The system’s adeptness at deciphering natural language instructions and materializing these into precise spatial arrangements has been demonstrated across various contexts, laying the groundwork for future advancements in multi-robot coordination. While our study focused on static patterns and centralized control, the insights gained provide a stepping stone for exploring dynamic pattern formations. ZeroCAP’s performance in context-aware task execution signal a promising direction for the integration of AI and robotics, propelling us towards more intelligent, adaptable, and efficient multi-robot applications in diverse real-world settings.

### v2

@@
 This not only validates the system’s capability to interpret and implement intricate context-driven tasks but also underscores its adaptability and effectiveness across varied environments and scenarios.
+The experimental videos and additional information about this work can be found at https://sites.google.com/view/zerocap/home.
 I Introduction
-Pattern formation, a key idea in the organization of multi-robot systems [1], involves arranging robots in a specific way to reach a common objective.
-This skill is essential for various purposes such as environmental surveillance, precise farming, cooperative transportation, and emergency management.
-The significance of pattern formation lies in its capacity to utilize the combined actions of robots to carry out tasks that go beyond what individual robots can achieve, thus improving effectiveness, adaptability, and robustness.
-Central to this endeavor is not just the geometric arrangement of robots, but the evolution towards purposeful pattern formations—where the design and execution of patterns are intrinsically linked to the context and objects within the environment.
-This nuanced approach marks a significant departure from traditional methodologies, steering the field towards more interactive and object-centric tasks that can be applicable to real world scenarios as shown in Fig 1.
+Pattern formation, a key idea in the organization of multi-robot systems (MRS) [1], involves arranging robots in a specific way to reach a common objective.
+This ability is critical for applications such as environmental surveillance, precision farming, cooperative transportation, and emergency management.
+Its value lies in coordinating robots to perform tasks beyond the capabilities of individuals, enhancing effectiveness, adaptability, and robustness.
+In this paper, we go beyond traditional geometric arrangements, focusing on purposeful pattern formations-where patterns are closely tied to user provided contexts and environments.
+This approach marks an evolution in MRS, steering the field towards more interactive and object-centric tasks that can be applicable to real world scenarios as shown in Fig 1.
 Despite its significance, traditional approaches to pattern formation in robotics have been predominantly constrained by model/rule-based methods and machine learning paradigms.
-These methodologies, while foundational, have their limitations: model rule-based strategies [1, 2, 3, 4], are often tailored to specific scenarios/patterns and lack the flexibility to adapt to new or unmodeled patterns.
-Reinforcement learning (RL) approaches, on the other hand, require training time and extensive reward design, making them impractical for scenarios requiring immediate deployment.
-Moreover, machine learning (ML) based methods [5, 6, 7] demand large datasets of annotated or labeled data for training, a requirement that poses substantial challenges in dynamic or novel environments.
-These limitations across all the methodologies are particularly pronounced in scenarios where pattern formations must be conditioned with respect to specific objects of interest, underscoring the need for a more adaptable and intuitive methodology.
+These methodologies, while foundational, have their limitations: model/rule-based strategies [1, 2, 3, 4] are often scenario-specific and lack adaptability to new patterns, while reinforcement learning (RL) requires significant training and reward design, making it impractical for immediate deployment.
+Moreover, machine learning (ML) methods [5, 6, 7] rely on large, annotated datasets, which are challenging in dynamic environments.
+These limitations are especially evident when pattern formation must be conditioned on user-provided natural language inputs referring to specific objects, emphasizing the need for a more flexible and intuitive approach.
 Figure 1: A conceptual image illustrating the proposed Zero-Shot multi-robot Context Aware Pattern (ZeroCAP) formation system, beginning with an initial environment where a car is parked incorrectly.
 Upon receiving the command from human operator to “Surround the incorrectly parked car at the corners!”, the system identifies the target vehicle and autonomously positions the robot cones at strategic corner locations, showcasing the successful execution of a zero-shot context aware pattern formation task.
-Integrating language comprehension into robotic systems heralds considerable progress [8, 9], yet it poses unique challenges, especially in spatially-focused tasks such as pattern formation.
-While language models have shown a remarkable ability as general pattern machines [10] and offer a robust framework for parsing and generating human-like text, Vision Language Models (VLMs) falter in accurately localizing spatial tasks, accentuating the need for techniques that adeptly marry abstract instruction understanding with their tangible implementation.
+Integrating language comprehension into robotic systems offers considerable progress [8, 9], but it also presents challenges in spatial tasks like pattern formation.
+While language models have shown a remarkable ability as general pattern machines [10] and offer a robust framework for parsing and generating human-like text, Vision Language Models (VLM) face challenges in precise spatial localization tasks due to limitations in spatial reasoning and dataset training [11, 12].
+This underscores the need for approaches that combine abstract instruction interpretation with accurate spatial execution by robots.
 Our contribution to this evolving field addresses the challenge previously identified by [3], focusing on the dynamic assignment of a variable number of robotic agents to specific target positions for intricate pattern formations.
-By combining vision-based processing with the sophisticated interpretative capabilities of Large Language Models (LLMs), we introduce a Zero-Shot multi-robot Context Aware Pattern (ZeroCAP) formation system that facilitates zero-shot, language, and context-conditioned pattern formation.
+By combining vision-based processing with the sophisticated interpretative capabilities of Large Language Models (LLMs), we introduce a Zero-Shot multi-robot Context Aware Pattern (ZeroCAP) formation system that facilitates zero-shot, language, and context-conditioned pattern formation without any prior training or data.
 It processes both images and natural language instructions to identify objects of interest and generate precise coordinates for robot deployment, enabling immediate and purposeful pattern formation directly influenced by user instructions.
-We distinguish our approach by categorizing pattern formations into three purposeful types: 1) General pattern formation, where robots arrange themselves around the object of interest based on user specification; 2) Infill pattern formation, where robots occupy the interior space of a region based on user specification; and 3) Caging pattern formation, creating a boundary around the object.
-This categorization increases the flexibility of deployment and signifies a shift from standard geometric patterns to contextually relevant formations tailored to the specific user context.
-Introducing purposeful pattern formation, our system adapts to environmental cues through natural language instructions, demonstrating language and context-driven pattern formation’s viability.
-The following sections will explore the technical challenges addressed, methodologies applied, and the significant impact our work has on advancing robotics towards intuitive, flexible, and context-aware multi-robot systems.
+The main contributions of this paper are as follows:
+  • ZeroCAP, a novel zero-shot multi-robot pattern formation framework that integrates LLMs with vision-based processing to translate user natural language inputs into context-aware, purposeful pattern formations, moving beyond traditional geometric arrangements.
+  • Implementation and evaluation of the framework in both simulation and real-world environments, validating the system’s ability to perform complex pattern formations with varied objects and natural language inputs.
+The following sections will explore the challenges tackled, methodologies employed, and the broader implications of our work for advancing intuitive, flexible, and context-aware MRS.
 II Related Works
 Pattern Formation.
-Multi-robot pattern formation research has explored centralized control [11], [12] and distributed approaches [1], [13].
-Centralized systems face scalability and adaptability challenges, while distributed methods, including consensus-based [14], leader-follower [15], and bio-inspired [16], improve scalability but lack robustness due to pattern-specific, rule-based strategies.
-Advancements to the feild include potential field-based methods[17], [4] and artistic formations [3], plus novel algorithms such as event-triggered mechanisms [2].
-However, model/rule-based strategies typically are designed for specific scenarios, lacking the adaptability required for new or unseen patterns or even handling user specific instructions.
-Conversely, learning-based strategies like Q-learning [7], [6]while offering more generalizability are often time intensive with their training and are limited to the fixed patterns that the system was trained for.
-Despite the above progress over the years, significant challenges persist in creating flexible solutions that quickly adapt to contextual user preferences and situations without heavy pre-configuration, training, or extensive information exchange.
-These issues highlight a broader need in the feild for moving from traditional geometric patterns to contextually relevant formations tailored to context specific user tasks and environments.
-Our approach diverges from these traditional paradigms by adopting zero-shot learning using a language model for robotics, sidestepping the limitations of both rule-based and machine learning methodologies.
-By leveraging natural language instructions, our system achieves a flexible, intuitive means of directing multi-robot pattern formation, capable of adaptation to novel scenarios without the need for extensive dataset training or predefined rules.
-This represents an advancement over existing methods centralized or decentralized.
+Multi-robot pattern formation has been explored through both centralized [13], [14] and distributed methods [1], [15].
+Centralized systems, while efficient in task coordination, can suffer from single points of failure and may be less adaptable in highly dynamic environments.
+Distributed approaches, such as consensus-based [16], leader-follower [17], and bio-inspired methods [18], improve adaptability and eliminate single points of failure but often rely on rule-based strategies that lack flexibility when applied to diverse or complex tasks.
+Advancements like potential field methods [19], [4], and event-triggered mechanisms [2] have been proposed, yet rule-based models still lack adaptability to novel or unseen patterns, while approaches like Mora et al.
+[3], emphasize visually appealing, artistic robot movements.
+Learning-based strategies like Q-learning [7], [6], while more generalizable, require extensive training and are limited to fixed patterns.
+Despite progress, challenges remain in developing flexible solutions that quickly adapt to user-specific contexts without pre-configuration or heavy training, moving beyond traditional geometric patterns to more contextually relevant formations.
+Our approach bypasses these limitations by employing zero-shot learning with a language model, enabling flexible, intuitive multi-robot pattern formation.
+Leveraging natural language instructions, our system adapts to novel static scenarios without extensive training or predefined rules, marking a significant improvement over existing methods.
 LLM for Robotics.
@@
 In the realm of multi-agent systems, LLMs have been instrumental in fostering agent cooperation, equipping individual agents with the capabilities to improve interaction dynamics and collaborative efficacy [24, 25, 26].
-Despite their growing utility in robotics, none of these noteable methodologies cover pattern formation among multi-robot systems.
-Vision Language Models, while useful in many applications involving robot navigation[25] and high-level action control[26, 27], are limited by their spatial reasoning capabilities [28], impacting their effectiveness in robotics applications that require precise spatial coordination.
-Our research leverages VLMs and LLMs for multi-robot pattern formation, addressing the spatial task limitations of VLMs and pioneering a linguistically-driven approach to complex robotic coordination aided by representing shapes using a node graph.
+Despite their growing utility in robotics, none of these noteable methodologies cover pattern formation among MRS.
+VLM, while useful in many applications involving robot navigation[27] and high-level action control[28, 29], are limited by their spatial reasoning capabilities [12], impacting their effectiveness in robotics applications that require precise spatial coordination.
+Our research leverages VLMs and LLMs for multi-robot pattern formation, addressing the spatial task limitations of VLMs and pioneering a natural language-driven approach to complex robotic coordination aided by representing shapes using a node graph.
 III Problem Description
-Our research delves into the nuanced task of deploying robotic agents to form patterns around or inside a designated object of interest $O$, located within an environment, $e$.
-This environment is depicted through an image ($I_{E}$), paired with a natural language instruction ($Ins$) that outlines the desired pattern formation task.
-Operating under a centralized control system, robots, simplified to points of mass, are coordinated to ensure precise pattern formation in the context of $O$ and $Ins$.
+This research focuses on the task of deploying robotic agents to form patterns around or inside a designated object of interest ($O$) within an environment ($e$), depicted through an image ($I_{E}$) and guided by a natural language instruction ($Ins$).
+We operate under a centralized control system where robots are simplified as point masses and coordinated to achieve precise pattern formation based on $O$ and $Ins$.
+This centralized approach is adopted to streamline task coordination by eliminating the need for each robot to independently process language and spatial data.
+It reduces computational overhead and avoids the significant costs of equipping each robot with its own VLM/LLM.
 The procedure begins with the identification and segmentation of $O$ from $I_{E}$, leading to a mask image that segregates $O$ from the rest of $e$.
@@
 This assignment facilitates the formation of the specified pattern around/inside $O$, in accordance with $Ins$.
+While the current formulation operates in 2D, based on the 2D input image $I_{E}$, it can be generalized to n-dimensions.
 Formally, this process can be encapsulated as a function $F$, where $F(I_{E},Ins)\rightarrow C_{xy}$, mapping the inputs to the deployment coordinates for each robot.
+This function represents the ZeroCAP system, which is detailed further in Sec.
+IV.
 Figure 2: An overview of the ZeroCAP system.
-It traces the workflow from the initial natural language instruction and input image of the environment through to the final deployment of robots, illustrating the sequence of processing stages—including context identification using Vision Language Model (VLM), object segmentation, shape description, and Large Language Model (LLM) coordination for precise robot placement in the environment.
+It traces the workflow from the initial natural language instruction and input image of the environment, to the final deployment of robots, illustrating the sequence of processing stages—including context identification using Vision Language Model (VLM), object segmentation, shape description, and Large Language Model (LLM) coordination for precise robot placement in the environment.
 Three key stages are highlighted and explained in Section IV.
 IV ZeroCAP Formation System
-This section delineates the comprehensive methodology of our ZeroCAP formation system, which introduces a novel approach for zero-shot multi-robot context aware pattern formation.
-Leveraging advanced VLMs, the methodology integrates linguistic instructions with visual perception, enabling precise spatial reasoning and object identification in multi-robot systems.
-Our approach harnesses the synergy of cutting-edge segmentation techniques, and utilizes standard computer vision algorithms for obtaining shape description.
-The decision-making process is empowered by a specialized LLM, which synthesizes the spatial description of object and linguistic context to generate target deployment coordinates for the robotic agents.
-The multi-robot control is centrally orchestrated, ensuring efficient navigation.
-The following subsections provide a detailed exposition of each stage, from initial input processing to the final deployment of robots.The overall structure of the proposed ZeroCAP framework is shown in Fig.
+This section outlines the methodology behind the ZeroCAP formation system.
+The system facilitates zero-shot, context-aware pattern formation for MRS by integrating LLMs with visual perception and analysis.
+The system, represented by the function $F$, takes two key inputs: $I_{E}$ and $Ins$.
+The VLM processes these inputs to identify $O$ and generate a pattern formation instruction ($PI$).
+The system then segments the object, extracts shape descriptors, and the LLM uses this spatial descriptors along with the instruction to generate precise deployment coordinates for the robots.
+Finally, under centralized control, the robots execute the task based on the generated coordinates.
+The overall structure of the ZeroCAP framework is depicted in Fig.
 2.
+Algorithm 1 outlines the execution pipeline of the system, from processing the inputs to deploying robots to the designated coordinates.
 IV-A Context Identification using VLM
-The initiation of our methodology engages a VLM to interpret an environmental image $I_{E}$ and a natural language instruction $Ins$.
-The architecture of this model excels at correlating visual stimuli with corresponding linguistic descriptors, emulating human-like proficiency in image-language tasks.
-Provided with $I_{E}$ and $Ins$, the VLM processes these inputs in tandem, producing two key outputs: the Object of Interest $O$ and the Pattern Formation Instruction $PI$.
-The Object of Interest $O$ is identified utilizing the VLM’s zero-shot detection capabilities, circumventing the need for prior object-specific training and ensuring versatility within varying environments.
-Conventional object detection methods may localize objects but lack the sophisticated reasoning to discern context as VLMs do.
-Given an instruction akin to that of Fig.
-3, standard natural language-aided object detectors [5] would falter in reasoning with which object is odd.
-In contrast, the advanced reasoning capabilities of the VLM, contextualized by the instruction, render it superior for identifying objects of interest within complex scenarios.
-This nuanced understanding of context allows the VLM not only to detect but also to reason about objects in a manner aligned with the given linguistic directives.
-Concurrently, the Pattern Formation Instruction $PI$ is extrapolated from the natural language directive enriched by contextual nuances in $I_{E}$.
+This stage, represented as box ‘Sec IV.
+A’ in Fig.
+2, begins with a VLM interpreting the environmental image $I_{E}$ and $Ins$.
+The VLM is proficient at correlating visual and linguistic inputs, emulating human-like proficiency in image-language tasks.
+This stage produces two key outputs: $O$ and $PI$.
+$O$ is detected using the VLM’s zero-shot capabilities, eliminating the need for prior object-specific training and allowing versatility in diverse environments.
+Unlike conventional object detection, which may localize objects but lacks contextual reasoning, the VLM can reason about the context and identify complex instructions [12].
+Taking the case of Fig.
+3, standard natural language-aided object detectors [5] struggle to reason about which object is odd, whereas the VLM’s advanced reasoning, guided by the user instruction, makes it more effective at identifying objects in complex scenarios.
+This advanced reasoning enables the VLM to align object identification with the given language directives.
+Simultaneously, the $PI$ is derived from the natural language directive, enriched by the context in $I_{E}$.
 Here, the VLM generates $PI$, inspired by chain-of-thought reasoning [31], to discern spatial and context-sensitive task directives.
-In Algorithm 1, Line 1 corresponds to estimating the context using the VLM.
-One notable strength of our system is our approach to mitigating the common issue of precise spatial localization, a challenge in standard VLMs.
-The inherent limitations of VLMs in tasks that require exact positional outputs make it impractical to use these models to directly provide the deployment coordinates $C_{xy}$.
-Instead, the VLM is optimized to identify $O$ with necessary precision and to extract a context-rich $PI$.
-These are indispensable for subsequent segmentation and precise pattern formation stages conducted by more specialized subsystems.
-By honing in on general features and attributes discernible in the environment and the language input, the VLM identifies $O$ with the necessary precision for subsequent segmentation and pattern formation stages.
-The VLM’s output, consisting of $O$ and $PI$, thus propels the ZeroCAP system into the segmentation phase, where $O$ is visually distinguished within $I_{E}$, and $PI$ informs the strategic pattern formation instruction to be implemented by the robotic entities.
+Algorithm 1, Line 1 corresponds to estimating the context using the VLM.
+One key strength of our system is addressing the challenge of precise spatial localization, a known limitation of standard VLMs.
+VLMs are not well-suited to directly generate $C_{xy}$, so instead, they focus on accurately identifying the $O$ and extracting a context-rich $PI$.
+These outputs are crucial for subsequent segmentation and pattern formation, handled by more specialized subsystems below.
+By focusing on general features and context from the environment and language input, the VLM ensures accurate identification of $O$ for further processing.
+The VLM’s output, consisting of $O$ and $PI$, thus propels the ZeroCAP system into the segmentation phase, where $O$ is visually distinguished within $I_{E}$, and $PI$ informs the strategic pattern formation instruction to be implemented.
 Figure 3: An illustration of the VLM processing a natural language instruction to identify the object of interest and generate a pattern formation instruction, which guides the deployment of robots in a specified arrangement around the object.
@@
 IV-B Object Segmentation and Shape Description
-Following the identification of $O$ and the acquisition of the $PI$ from the VLM, the system transitions into the crucial stages of segmentation and shape description.
-These stages leverage sophisticated language conditioned segmentation techniques [5, 31], to perform image segmentation, followed by shape description stage using standard computer vision methods.
+Following the identification of $O$ and acquiring the $PI$ from the VLM, the system moves to segmentation and shape description represented as box ‘Sec IV.
+B’ in Fig.
+2.
+These stages utilize language-conditioned segmentation [5, 32], and standard computer vision methods for shape analysis.
 IV-B1 Object Segmentation
-The segmentation process begins with language-prompt-based object segmentation, which employs LangSAM, an open-source language-conditioned segmentation model built upon segment-anything approach and the GroundingDINO detection model [31, 32].
-Provided with the Object of Interest $O$ and the Input Image $I_{E}$, these models generate a binary segmentation map in which the object is highlighted in white against a black background.
-Upon completion, the segmentation yields a Masked Image $I_{M}$ and an accuracy metric $\alpha$, compared against a predefined threshold $t$.
-If $\alpha$ meets or exceeds this threshold, the process proceeds, confirming the segmentation’s quality for subsequent tasks.
-This segmentation approach ensures the precise extraction of $O$ from $I_{E}$, encapsulated within $I_{M}$ for the next processing stage.
-Lines 2 to 4 of the Algorithm 1 expound the segmentation process.
+The segmentation process employs LangSAM, an open-source language-conditioned model based on the segment-anything approach and GroundingDINO [32, 33].
+Given $O$ and $I_{E}$, the model generates a binary segmentation map, highlighting the object in white against a black background.
+The segmentation outputs a Masked Image $I_{M}$ and an accuracy metric $\alpha$, which is a scalar value compared against a threshold $t$, a predefined scalar constant.
+If $\alpha$ exceeds $t$, the process continues, ensuring the segmentation quality is sufficient.
+This approach precisely extracts $O$ from $I_{E}$, encapsulated in $I_{M}$ for further processing (Algorithm 1, Lines 2-4).
 IV-B2 Shape Description
-With a high quality $I_{M}$ at hand, the ZeroCAP system advances to the shape description phase.
-Here, standard computer vision techniques preprocess $I_{M}$, including filtering and enhancement operations to sharpen feature delineation.
-Through OpenCV’s edge detection algorithms and contour detection methods, $I_{M}$ is analyzed to determine $O$ geometric form.
-This analysis yields the object’s edges $E$ and vertices $V$, forming a connected graph $G=(V,E)$ that mathematically represents the object’s structure.
-Algorithm 1 delineates this process in Lines 10 – 13, with a method call in Line 5.
-Integrating these segmentation techniques with precise shape description, the ZeroCAP system ensures accurate recognition, segmentation, and modeling of the Object of Interest.
-This forms a solid foundation for the ensuing methodology phase: LLM-based Pattern Former.
+With a quality $I_{M}$, the system proceeds to shape description.
+Using standard computer vision techniques, $I_{M}$ is preprocessed for edge and contour detection through OpenCV, resulting in the identification of edges $E$ and vertices $V$.
+These form a connected graph $G=(V,E)$, which mathematically represents $O$.
+The process is shown in Algorithm 1, Lines 9-12, initiated by Line 5.
+ZeroCAP thus integrates segmentation and shape description to ensure accurate object recognition and modeling, forming the basis for the LLM-based Pattern Former.
 Algorithm 1 ZeroCAP System Pipeline
@@
 8:$C_{xy}\leftarrow\textsc{LLM.GenerateCoordinates}(PI,V,E,R)$
-9:for each coordinate $(x_{i},y_{i})$ in $C_{xy}$ do
-10:   $r_{i}\leftarrow R[i]$
-11:   DeployRobots($r_{i},(x_{i},y_{i})$)
-12:procedure ShapeDescriptor($I_{M}$)
-13:   Preprocess the masked image $I_{M}$
-14:   Extract edges $E$ and vertices $V$
-15:   return $(V,E)$ from $I_{M}$
-16:procedure DeployRobots($r,(x,y)$)
-17:   Command robot $r$ to move to coordinates $(x,y)$
-18:   Verify that robot $r$ has reached $(x,y)$
+9:for each robot $r_{i}$ in $R$ and coordinate $(x_{i},y_{i})$ in $C_{xy}$ do
+10:   DeployRobots($r_{i},(x_{i},y_{i})$)
+11:procedure ShapeDescriptor($I_{M}$)
+12:   Preprocess the masked image $I_{M}$
+13:   Extract edges $E$ and vertices $V$
+14:   return $(V,E)$ from $I_{M}$
+15:procedure DeployRobots($r,(x,y)$)
+16:   Command robot $r$ to move to coordinates $(x,y)$
+17:   Verify that robot $r$ has reached $(x,y)$
 IV-C LLM-based Pattern Former
-Upon accurately delineating the geometry of $O$, the context-aware multi-robot deployment system advances to determining goal position coordinates using a LLM.
-This stage involves the LLM being provided with the number of available robots $R$, interpreting the Pattern Formation Instruction $PI$ alongside the object’s spatial characteristics, represented through edges $E$ and vertices $V$ to provide an output representing positions for robots.
-The LLM’s task is to integrate the $PI$ with the spatial data of $E$ and $V$ to determine the optimal pattern formation.
+After delineating the geometry of $O$, the system advances to generating goal position coordinates using a LLM represented as box ‘Sec IV.
+C’ in Fig.
+2.
+The LLM receives the number of available robots $R$, $PI$, and the spatial characteristics of the object, represented by its edges $E$ and vertices $V$, to compute the optimal robot deployment positions.
+Instead of using the entire Masked Image $I_{m}$, we opt for the more computationally efficient use of $E$ and $V$, significantly reducing token overhead for LLMs and enhancing real-time performance.
+This approach ensures faster processing while keeping computational costs low.
+The LLM integrates the $PI$ with the spatial data $E$ and $V$ to determine the optimal pattern formation.
 This involves comprehensive reasoning to align the robots according to $PI$, while considering the shape of the object and the available robotic agents $R$.
-In particular, our methodology eschews the direct use of the entire Masked Image $I_{m}$ as input to the LLM, a decision underscored by considerations of computational efficiency and token costs.
-Direct image is resource-intensive, whereas leveraging $E$ and $V$ markedly reduces computational load especially when accounting for tokens in the context of LLM’s.
-This choice highlights our system’s capacity for real-time processing and deployment efficiency while also remaining inexpensive in computation and practicality.
-The culmination of this process is the production of a coordinate list $C_{xy}$ as seen in Line 6 of Algorithm 1, which prescribes exact deployment coordinates for each robot.
+The output of this stage is a list of coordinates $C_{xy}$ specifying precise deployment locations for each robot, as outlined in Line 6 of Algorithm 1.
 IV-D Robot Deployment
-Upon generating the deployment coordinates $C_{xy}$ with the LLM, the ZeroCAP system initiates robot deployment, translating algorithmic output into coordinated physical actions.
-Each robot $r_{i},i\in N$ is directed toward its assigned location $(x_{i},y_{i})$ under centralized control, ensuring precise navigation.
-Upon arrival, the system verifies each robot’s position against $C_{xy}$, adjusting as necessary to rectify any positional inaccuracies, thereby ensuring the formation aligns perfectly with the pattern instructions.
-This is delineated in Lines 7 to 9 and 14 to 16 in Algorithm 1.
-The successful deployment of robots in the context of the Object of Interest $O$ marks the completion of the pattern formation task, embodying the ZeroCAP system’s capability to execute context-driven, instruction-based patterns through autonomous robotic agents.
+Once the deployment coordinates $C_{xy}$ are generated, the ZeroCAP system initiates robot deployment, translating these coordinates into coordinated actions.
+Each robot $r_{i},i\in N$ is directed toward its assigned location $(x_{i},y_{i})$.
+Upon arrival, the system verifies each robot’s position against $C_{xy}$, making adjustments as needed to correct any deviations, ensuring the formation aligns with the pattern instructions.
+This is shown in Lines 7 to 8 and 13 to 15 in Algorithm 1.
+The successful deployment of robots around the $O$ completes the pattern formation task, demonstrating ZeroCAP’s ability to execute context-aware, instruction-based patterns, without the need for any prior training.
 V Experiment
-(a) General pattern formation tasks: robots arrange themselves around the object of interest based on user specification
-(b) Infill pattern formation tasks: robots occupy the interior space of a region based on user specification
-(c) Caging formation tasks: creating a cage like boundary around the object
-(d) Application oriented task: surrounding the incorrectly parked car with four robots.
-Figure 4: Illustrations of the three types of pattern formation tasks and a task demonstration of Fig 1, showcasing the system’s adaptability to various instruction types and object configurations.
-In all tasks, robots are represented as black points initially distributed at the bottom of the scene in the initial environment, transitioning to form user-directed patterns in the final environment.
-This illustrates the dynamic capability of our system to interpret instructions and spatially reorganize the robots accordingly.
+Figure 4: Illustrations of five real-world tasks performed by the ZeroCAP system, showcasing different pattern formations: (a) and (b) illustrate general pattern formations; (c) and (d) depict infill pattern formations; (e) demonstrates a caging task.
+Each task is executed based on natural language instructions provided at the top of each subfigure.
+Blue boundaries around the robots indicate their initial positions, while green boundaries represent their final positions.
 Please zoom in on the images for details.
-Our experimental evaluation investigates the ZeroCAP system’s zero-shot pattern formation capabilities across ten tasks as shown in Fig.
-4, delineated into three categories as prefaced in Section I: general pattern formations as shown in subfigure 4(a), infill pattern formations as shown in subfigure 4(b), and caging formation tasks as shown in subfigure 4(c) and the task referred to in Fig 1 as shown in subfigure 4(d).
-Tasks are diversified by object setup to evaluate the system’s adaptability and discernment capabilities under varying environmental complexities.
-In the Single Object setup (ENV a, b, d, g), the environment contains a solitary object, simplifying the identification process.
-The Multi-Object setup (ENV c, e, h) introduces additional objects as distractors, challenging the system to isolate the target object based on explicit user instructions.
-Lastly, the Hidden Object setup (ENV f, i, park) presents a more complex scenario where the natural language instruction implies the object of interest without explicit mention, compelling the system to employ reasoning to deduce the correct object.
+We evaluate the ZeroCAP system’s zero-shot pattern formation capabilities across five real-world tasks, as shown in Fig.
+4 and ten tasks performed in simulated environments.
+These tasks are categorized into three types: 1) General pattern formation, where robots arrange around the object of interest; 2) Infill pattern formation, where robots occupy the interior of a region; and 3) Caging formation, where robots create a boundary around the object.
+This categorization shifts from standard geometric patterns to contextually relevant formations, driven by natural language instructions and environmental cues, demonstrating the system’s flexibility.
+The tasks also vary by object setup to test adaptability.
+In the Single Object setup as shown in Fig.
+4, a single object simplifies identification.
+The Multi-Object setup, shown in Fig.
+5 adds distractors, requiring the system to isolate the target from several options based on explicit instructions.
+In the Hidden Object setup, shown in Fig.
+5 and Fig.
+6, the target is implied rather than stated, requiring the system to reason based on contextual clues.
 V-A Baseline Comparisons
-Our system’s performance was benchmarked against three baselines.
-The first is an RL-based simplified approach with hand-crafted rewards, serving primarily as a pattern-formation benchmark without natural language comprehension.
-The second and third baselines use GPT4 Vision [33] and llava-v1.6 [34], respectively, relying solely on vision-language models to complete the tasks.
-The LLMs employed in our experiments are GPT-4 [35], llama70b [36], and Claude [37], contrasting their performance against each other and the baselines.
-For ablation studies, the default configuration of the ZeroCAP system utilizes GPT-4 as the primary LLM.
+The system is benchmarked against three baselines.
+The first is an RL-based approach with hand-crafted rewards, focusing on pattern formation without natural language comprehension.
+The second and third baselines utilize GPT-4 Vision [34] and LLava-v1.6 [35], relying solely on vision-language models to complete tasks.
+The LLMs used in our experiments include GPT-4 [36], llama70b [37], and Claude [38], allowing for performance comparison across models and baselines.
+In ablation studies, GPT-4 is the default LLM configuration for ZeroCAP.
 V-B Environment
-Experiments were conducted in a 2D simulation using Pygame, with robots as point-mass entities capable of precise movements to target coordinates.
-The setup includes multi-object tasks with infill and caging requirements, omitting obstacle avoidance behaviors to focus on the pattern formation tasks.
-A subset of tasks was also executed on a real testbed, featuring Hamster mobile robots [38] under an overhead camera setup.
+Experiments were conducted in both real-world settings and in 2D simulations.
+Real-world experiments used a testbed with Hamster mobile robots under an overhead camera [39], with the environment simplified into a grid-based image to ensure consistency in point-mass robot deployment.
 Fig.
-5 showcases two tasks executed in a real-world setting, serving as a tangible proof of concept that corroborates our simulation results.
-To facilitate these tasks, the environment is simplified into a grid-based image, with each grid box approximately matching the size of a robot.
-This adaptation allows the robots to retain their point mass entity characterization within the practical constraints of the real-world experiments, ensuring consistency in their deployment and movement.
-For demonstrations of simulator and real-world tasks, please refer to the supplementary video.
+4 highlights the real-world tasks, demonstrating the practical viability of our approach.
+Additionally, 2D simulations were carried out using the Pygame environment, as shown in Fig.
+5, where robots were modeled as point-mass entities and moved precisely to target coordinates.
+The simulation setup included multi-object tasks with infill and caging requirements, excluding obstacle avoidance to focus solely on pattern formation.
+These simulation experiments served to rigorously validate the proposed methods across varied conditions and object setups.
+For more detailed demonstrations of the system under various setups and task categories, please refer to the supplementary video.
 V-C Evaluation Metrics
-We employed two main evaluation metrics: Success Rates (SR) and Goal Condition Recall (GCR).
-SR indicates whether the robots formed the pattern successfully, verified by handcrafted ground truth labels, while GCR quantifies the proportion of robots correctly positioned within the pattern.
-A GCR of 1 implies a perfect match with the ground truth and, consequently, an SR of 1.
-Each task underwent 10 trials, with averages reported for SR and GCR, chosen to account for the inherent variability in LLM performance.
+We used two evaluation metrics: Success Rate (SR) and Goal Condition Recall (GCR).
+SR indicates if the robots successfully formed the pattern by comparing final positions to handcrafted ground truth labels, with success requiring all positions to match.
+GCR measures the proportion of correctly positioned robots, with a GCR of 1 implying an SR of 1.
+Each task was run over 10 trials, and average SR and GCR were reported to account for LLM variability.
 VI Results and Analysis
-TABLE I: Comparative analysis of pattern formation methodologies across General, Infill, and Caging tasks, evaluating Success Rates (SR) and Goal Condition Recall (GCR) metrics across different methods.
+Table I summarizes the evaluation of the tested methodologies, highlighting ZeroCAP’s effectiveness over the baselines.
+GPT-4 Vision and LLava v1.6, are tested for direct VLM-only zero-shot pattern formation, both performed poorly.
+GPT-4 Vision failed to generate meaningful coordinates, revealing its limitations in spatial reasoning.
+LLava produced inaccurate or irrelevant positions.
+These results show that using VLMs alone for context-aware pattern formation is inadequate without enhanced spatial reasoning capabilities.
+The RL-based baseline showed mixed results, struggling with general and infill pattern formation due to varied user instructions requiring flexibility.
+RL’s reliance on predefined rewards limits generalizability; while effective in task-specific training, it underperforms with generalized reward structures.
+We chose a generalized reward approach to enhance scalability, as designing specific rewards for every new task is impractical.
+Consequently, RL performed poorly in general and infill tasks but excelled in caging tasks, where the consistent goal of forming a barrier aligned with the reward design.
+This highlights RL’s strength in structured tasks but its limitations in adapting to nuanced, context-driven user instructions.
+In contrast, the ZeroCAP system, leveraging LLMs like GPT-4, llama-2-70b, and Claude-3-opus, significantly outperformed the baselines.
+GPT-4-based ZeroCAP achieved the highest SR and GCR across all tasks, largely due to its superior language processing, allowing for better interpretation of complex instructions.
+The ZeroCAP system’s architecture, which decouples spatial reasoning from the VLM and utilizes edge representations, mitigated GPT-4 Vision’s limitations in direct spatial localization.
+Llama-2-70b performed moderately well but showed inconsistencies, particularly in tasks with hidden object scenarios requiring complex spatial reasoning.
+Claude-3-opus, while slightly less accurate than GPT-4, demonstrated more consistency across different task types.
+Figure 5: Two simulation scenarios demonstrating ZeroCAP’s ability to execute context-aware pattern formations.
+Sim1 (left) depicts a caging task in a hidden object setup, where the object of interest is not directly specified and must be inferred.
+Sim2 (right) illustrates an infill task in a multi-object setup.
+Please zoom in on the images for details.
+Figure 6: Illustration of a general pattern formation task as shown in concept Fig.
+1, featuring a hidden object setup where the system must infer the wrongly parked car and deploy robots, represented by black points, to surround it.
+Task-wise, general pattern formation proved the most challenging, likely due to the abstract nature of the instructions, requiring nuanced interpretation.
+Infill tasks, with clearer objectives, achieved the highest SR and GCR, particularly with GPT-4, as they mainly involved placing robots within a specified region.
+Caging tasks, although more structured, still posed difficulties, but ZeroCAP managed consistent barrier formations effectively.
+TABLE I: Comparative analysis of pattern formation methods for General, Infill, and Caging tasks, evaluating Success Rates (SR) and Goal Condition Recall (GCR) metrics across different methods.
+Total represents the combined performance across all task categories.
+Bold numbers indicate the best performance among the tested methods.
 Methods General Infill Caging Total
@@
 GPT-4 Vision (Only) 0 0 0 0 0 0 0 0
-LLaVA v1.6 (Only) 0 0 0.13 0.1 0 0.07 0 0.07
+LLaVA v1.6 (Only) 0 0 0 0.15 0 0.07 0 0.22
 RL Baseline 0.35 0.53 0.28 0.41 0.88 0.9 0.5 0.61
 ZeroCAP (GPT-4) 0.71 0.85 0.86 0.89 0.74 0.82 0.77 0.85
-ZeroCAP (lama-2-70b) 0.2 0.47 0.39 0.71 0.32 0.67 0.3 0.61
-ZeroCAP (claude-3-opus) 0.57 0.78 0.72 0.88 0.64 0.73 0.64 0.79
-Table I shows a comprehensive evaluation of different methodologies tested in this paper, revealing the effectiveness of the ZeroCAP system in comparison to baselines.
-The initial baselines, GPT-4 Vision and LLava v1.6, were employed to evaluate the feasibility of using VLMs directly for zero-shot context-aware pattern formation.
-The performance of these baselines was significantly poor.
-GPT-4 Vision generally failed to generate any coordinates, reflecting its known limitations with spatial localization tasks.
-LLava, while capable of producing coordinates, generated positions that were largely inaccurate and contextually irrelevant.
-These outcomes effectively demonstrate that direct application of VLMs for context-aware pattern formation is non-trivial, and without additional processing or reasoning mechanisms, VLMs fall short of providing viable solutions.
-The RL baseline exhibited a mixed performance across the task spectrum.
-It notably struggled with the general pattern formation tasks, where user specifications varied greatly and required a high degree of interpretation and flexibility.
-The RL method’s rigidity and lack of language comprehension rendered it unable to adapt to the nuanced demands of these tasks.
-On the other hand, the RL approach showed considerable success in Caging tasks, aligning with the more straightforward objective of these tasks where the overarching goal is consistently to form a barrier around the object.
-Figure 5: Two real-world scenarios demonstrating ZeroCAP system’s ability to execute context-aware pattern formations.
-The ZeroCAP methods, incorporating LLMs like GPT-4, llama-2-70b, and claude-3-opus, present a contrasting picture.
-The GPT-4 based ZeroCAP achieves the highest overall SR and GCR across all types of tasks, indicating its superior pattern recognition and execution in zero-shot conditions.
-This suggests that GPT-4’s advanced natural language processing abilities are effectively leveraged by the ZeroCAP system to interpret instructions and generate appropriate robotic behaviors.
-The performance of ZeroCAP using llama-2-70b and claude-3-opus varies.
-While the llama-2-70b variant has some success, its inconsistency across tasks points to possible limitations in its ability to handle complex spatial reasoning as compared to GPT-4.
-Claude-3-opus shows a relatively more consistent performance.
-Across different pattern formation types, ZeroCAP methods behave in a manner that suggests certain patterns may be more conducive to LLM-driven execution.
-The General pattern formation tasks appear to be the most challenging, possibly due to the abstract nature of the instructions and the need for more nuanced interpretation.
-Infill tasks, which involve a more defined area and clear objectives being in most cases to deploy robots inside a region, show the highest SR and GCR for GPT-4, indicating a strong alignment between the method’s capabilities and the task requirements.
-Caging tasks, while more structured, still present a challenge, but ZeroCAP methods demonstrate reasonable effectiveness.
+ZeroCAP (llama-2-70b) 0.39 0.47 0.55 0.71 0.62 0.74 0.52 0.64
+ZeroCAP (claude-3-opus) 0.63 0.78 0.79 0.88 0.69 0.79 0.70 0.81
 VI-A Ablation: Segmentation Methods
-The ablation study presented in Table II examines the efficacy of CLIPSeg [5] and LangSAM as object segmentation methods within the context of language-prompted object identification.
-These methods, while capable of segmenting objects based on language instructions, do not possess the same reasoning abilities provided by full-fledged VLMs.
-The table compares the performance of these segmenters used in isolation (denoted by ‘Only’) to their performance when integrated into the ZeroCAP framework.
+Table II presents the results of an ablation study comparing the performance of CLIPSeg [5] and LangSAM as standalone segmentation methods and their integration within the ZeroCAP framework.
+While CLIPSeg and LangSAM achieve modest success rates (0.51 and 0.43, respectively) when used alone, their performance is limited by a lack of spatial reasoning and deeper contextual understanding, particularly in tasks requiring hidden object identification.
 TABLE II: Ablation studies comparing the success rates (SR) of standalone segmentation methods and their integration within the ZeroCAP system.
@@
 ZeroCAP with LangSAM 0.77
-It is evident that using CLIPSeg and LangSAM alone for object identification and segmentation results in modest success rates of 0.51 and 0.43, respectively.
-This suggests that while they can interpret natural language to some extent and perform segmentation, the lack of deeper contextual understanding and spatial reasoning affects their overall effectiveness in pattern formation tasks.
-This is especially a factor in case of hidden object task scenarios.
-When these segmentation methods are incorporated into the ZeroCAP system, the success rates improve significantly to 0.78 for CLIPSeg and 0.77 for LangSAM.
-The integration of these segmenters within ZeroCAP likely benefits from additional contextual processing and spatial reasoning capabilities provided by the LLMs in the system, leading to more successful task outcomes.
+When integrated into ZeroCAP, success rates rise to 0.78 and 0.77, respectively.
+This improvement demonstrates the added value of incorporating VLMs into the system, providing the necessary contextual reasoning and spatial capabilities to enhance overall performance.
 VI-B Ablation: Shape Descriptors
-The ablation study presented in Table III evaluates the influence of different shape descriptors on the performance of the ZeroCAP system.
-‘Increase in Tokens’ indicates the relative percentage increase in the number of tokens used by the LLM compared to using only ‘Edges’ as the shape descriptor, which serves as the baseline in this context.
+The ablation study presented in Table III examines the impact of different shape descriptors on ZeroCAP’s performance.
+‘Increase in Tokens’ reflects the percentage increase in token usage compared to using only ‘Edges’ as the shape descriptor.
 TABLE III: Ablation studies on the impact of different shape descriptors on Success Rates (SR) and their relative increase in tokens required for LLM processing.
@@
 Binary Image Matrix 0.71 400%
-Analyzing the data, it is evident that using only ‘Edges’ as a shape descriptor achieves the highest SR of 0.77.
-This suggests that ‘Edges’ alone provide ample information for the LLM to accurately generate the desired patterns without requiring additional token input.
-In contrast, when ‘Edges and Vertices’ are combined, there is a modest increase in tokens by 2%, but the SR significantly decreases to 0.48.
-This decrease could imply that including vertices introduces a level of complexity that the LLM does not effectively translate into better pattern formation.
-Moreover, vertices as a descriptor could potentially be misleading for non-polygonal shapes where the definition of vertices is less clear, potentially leading to inaccurate pattern formation by the LLM.
-The binary image matrix descriptor is obtained by converting the segmented object mask into a binary matrix, which is then resized to a 100x100 grid to fit within the token constraints of the LLM.
-This matrix serves as a input for the LLM, representing the shape in a form that can be directly processed.
-The LLM utilizes this grid-based representation to understand the spatial arrangement and generate target coordinates for the robots.
-The binary image matrix descriptor, incurs a 400% increase in token usage.
-This upsurge poses practicality concerns due to token limits and practical costs.
-Additionally, resizing may obscure vital shape details, affecting pattern accuracy.
-Despite a reasonable SR of 0.71, the trade-off between detail richness and token efficiency is crucial for the ZeroCAP system’s viability.
-VII Discussion
-The research presented in this paper marks a significant stride in the domain of multi-robot systems, particularly in the application of LLMs for zero-shot context aware pattern formation.
-Nonetheless, the current scope is confined to static patterns, leaving the exploration of dynamic patterns - those that evolve over time in response to environmental changes or ongoing tasks - as a promising avenue for future research.
-Dynamic pattern formation would not only extend the flexibility and adaptability of multi-robot systems but also increase their applicability in real-world scenarios such as search and rescue operations or environmental monitoring, where conditions are continually shifting.
-Our current method’s reliance on centralized processing poses potential bottlenecks and failure points in scenarios requiring quick robot adaptability.
-Exploring decentralized or hybrid control systems could improve robustness and scalability by merging centralized planning with decentralized execution.
-Additionally, incorporating sensory feedback mechanisms may enhance decision-making, maintaining system effectiveness amid real-time environmental changes.
-Our research lays a foundation for future advancements in robotic pattern formation and practical applications, extending from complex coordination tasks in surveillance, logistics, and warehousing to context-aware strategies.
-Progressing beyond current limitations and enhancing our foundational framework is vital for fully realizing robotics’ capabilities in intricate, context-sensitive environments.
-VIII Conclusion
-Our research has successfully introduced ZeroCAP, an innovative framework that harmonizes the advanced linguistic comprehension of large language models with the physical capabilities of robotic systems to achieve zero-shot pattern formation.
-The system’s adeptness at deciphering natural language instructions and materializing these into precise spatial arrangements has been demonstrated across various contexts, laying the groundwork for future advancements in multi-robot coordination.
-While our study focused on static patterns and centralized control, the insights gained provide a stepping stone for exploring dynamic pattern formations.
-ZeroCAP’s performance in context-aware task execution signal a promising direction for the integration of AI and robotics, propelling us towards more intelligent, adaptable, and efficient multi-robot applications in diverse real-world settings.
+Results show that using ‘Edges’ alone yields the highest SR (0.77), demonstrating that this descriptor provides sufficient information for the LLM without increasing token load.
+Adding ‘Vertices’ results in only a 2% token increase, but the SR drops to 0.48, indicating that the complexity introduced by vertices can hinder performance, especially for non-polygonal shapes where vertex definitions are less clear.
+The binary image matrix, created by converting and resizing the object mask to a $100\times 100$ grid, increases token usage by 400%, making it less practical.
+While it achieves a decent SR of 0.71, the token cost and potential loss of important shape features during resizing highlight the trade-off between detail and efficiency in ZeroCAP’s processing.
+Overall, the results confirm that VLMs alone are insufficient for zero-shot context-aware pattern formation, underscoring the need for ZeroCAP.
+ZeroCAP outperforms baselines, particularly in tasks requiring nuanced reasoning.
+Ablation studies show that edges as shape descriptors yield the best success with minimal token load, highlighting ZeroCAP’s strength in integrating LLMs for robust, zero-shot pattern formation and setting a foundation for future multi-robot coordination research.
+VII Conclusion
+This work presents ZeroCAP, a framework integrating LLMs with MRS for zero-shot, context-aware pattern formation.
+By translating natural language instructions into precise spatial arrangements, ZeroCAP overcomes VLM spatial localization limitations by decoupling spatial reasoning from visual processing, using edge and vertex representations for more accurate pattern formation.
+Currently, due to VLMs limitations in processing 3D data, being trained mostly on 2D images [11], this work is constrained to a 2D framework to ensure robustness.
+As VLMs evolve, future exploration could extend to 3D.
+The system is deployable with a map or overhead camera and could be expanded using multiple cameras in the future for wider scale and coverage.
+While our focus has been on static formations, dynamic formations that adapt to real-time changes hold significant promise.
+Further investigation of complex environments with partial occlusions could also improve robustness.
+This early-stage research lays the groundwork for more intuitive robotic coordination with implications in applications such as surveillance and logistics.
+Acknowledgement
+We would like to thank Dr. Tamzidul Mina at Sandia National Laboratories for his valuable feedback, which significantly enhanced the clarity and quality of this paper.

### v3

@@
 In this paper, we go beyond traditional geometric arrangements, focusing on purposeful pattern formations-where patterns are closely tied to user provided contexts and environments.
-This approach marks an evolution in MRS, steering the field towards more interactive and object-centric tasks that can be applicable to real world scenarios as shown in Fig 1.
+This approach marks an evolution in MRS, steering the field towards more interactive and object-centric tasks that can be applicable to real world scenarios as shown in Fig.
+1.
 Despite its significance, traditional approaches to pattern formation in robotics have been predominantly constrained by model/rule-based methods and machine learning paradigms.
@@
 V-A Baseline Comparisons
-The system is benchmarked against three baselines.
-The first is an RL-based approach with hand-crafted rewards, focusing on pattern formation without natural language comprehension.
-The second and third baselines utilize GPT-4 Vision [34] and LLava-v1.6 [35], relying solely on vision-language models to complete tasks.
-The LLMs used in our experiments include GPT-4 [36], llama70b [37], and Claude [38], allowing for performance comparison across models and baselines.
-In ablation studies, GPT-4 is the default LLM configuration for ZeroCAP.
+The system is benchmarked against three baselines: an RL-based approach with hand-crafted rewards for pattern formation—where Caging rewarded enclosing a target, Infill incentivized structured spacing inside a region, and General used a proximity-based reward for clustering robots near a reference point—and two vision-language models, GPT-4 Vision [34] and LLava-v1.6 [35].
+Experiments employ GPT-4 [36], LLaMA-70B [37], and Claude [38], with GPT-4 serving as the default in ablation studies.
+Unlike traditional rule-based pattern formation methods, ZeroCAP interprets natural language instructions to enable context-aware formations, rendering direct comparisons with classical approaches infeasible due to their lack of user intent integration.
 V-B Environment
@@
 1, featuring a hidden object setup where the system must infer the wrongly parked car and deploy robots, represented by black points, to surround it.
-Task-wise, general pattern formation proved the most challenging, likely due to the abstract nature of the instructions, requiring nuanced interpretation.
-Infill tasks, with clearer objectives, achieved the highest SR and GCR, particularly with GPT-4, as they mainly involved placing robots within a specified region.
-Caging tasks, although more structured, still posed difficulties, but ZeroCAP managed consistent barrier formations effectively.
+Task-wise, general pattern formation was the most challenging due to language ambiguity and the lack of structured references, leading to misinterpretations in robot placements.
+Infill tasks achieved the highest SR and GCR, though occasional failures occurred from uneven spacing inside the region.
+Caging tasks, while more structured, encountered difficulties with precise localization, often resulting in incomplete barriers due to gaps in robot placement, but ZeroCAP managed consistent barrier formations effectively.
 TABLE I: Comparative analysis of pattern formation methods for General, Infill, and Caging tasks, evaluating Success Rates (SR) and Goal Condition Recall (GCR) metrics across different methods.
@@
 By translating natural language instructions into precise spatial arrangements, ZeroCAP overcomes VLM spatial localization limitations by decoupling spatial reasoning from visual processing, using edge and vertex representations for more accurate pattern formation.
-Currently, due to VLMs limitations in processing 3D data, being trained mostly on 2D images [11], this work is constrained to a 2D framework to ensure robustness.
-As VLMs evolve, future exploration could extend to 3D.
-The system is deployable with a map or overhead camera and could be expanded using multiple cameras in the future for wider scale and coverage.
-While our focus has been on static formations, dynamic formations that adapt to real-time changes hold significant promise.
-Further investigation of complex environments with partial occlusions could also improve robustness.
-This early-stage research lays the groundwork for more intuitive robotic coordination with implications in applications such as surveillance and logistics.
+Currently, as VLMs are mostly trained on 2D images [11], ZeroCAP is confined to a 2D framework for robustness, though future developments may extend it to 3D.
+The system is deployable with a map or overhead camera and could be expanded using multiple cameras in the future for wider coverage.
+While our focus has been on static formations, dynamic, real-time adaptations in complex, occluded environments hold significant promise, paving the way for more intuitive robotic coordination in applications such as surveillance and logistics.
 Acknowledgement